# natural language processing; greenwashing detection; corporate climate-related greenwashing; climate pretrained models; climate topic detection; green claim detection; climate QA; climate-related deceptive techniques; enviromental performance prediction copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Natural language processing

### v1

@@ C.2. Thematic Analysis
(b) ROC-AUC scores on TCFD classification task (Sampson et al., 2022)
Source Model F1
Auzepy et al. (2023) BART+MNLI 56.68%
(c) F1-scores (macro average) on TCFD classification (Auzepy et al., 2023)
Table 13. Reported performances for the datasets on TFCD classification.
Source Model F1
Huang et al. (2020) FinBERT 89.6%
Huang et al. (2020) BERT 86.9%
Huang et al. (2020) NB 75.0%
Huang et al. (2020) SVM 75.0%
Huang et al. (2020) RF 78.3%
Huang et al. (2020) CNN 81.9%
Huang et al. (2020) LSTM 86.0%
(a) F1-scores (micro average) on ESG classification (Huang et al., 2020)
Source Model Accuracy
Lee and Kim (2023) Multiling. BERT 84.09% ±$\pm 0.092$
Lee and Kim (2023) KoBERT 85.98% ±$\pm 0.163$
Lee and Kim (2023) DistilKoBERT 84.89% ±$\pm 0.081$
Lee and Kim (2023) KB-ALBERT 85.62% ±$\pm 0.026$
Lee and Kim (2023) KLUE (base) 86.05% ±$\pm 0.125$
Lee and Kim (2023) KLUE (large) 86.66% ±$\pm 0.099$
(b) Accuracy on ESG classification (Lee and Kim, 2023). Accuracy reported as mean ±$\pm std$
Environment F1
Schimanski et al. (2023b) BERT 91.85% ±$\pm 1.25$
Schimanski et al. (2023b) RoBERTa 92.35% ±$\pm 2.29$
Schimanski et al. (2023b) DistilRoBERTa 90.97% ±$\pm 2.00$
Schimanski et al. (2023b) EnvRoBERTa 93.19% ±$\pm 1.40$
Schimanski et al. (2023b) EnvDistilRoBERTa 92.35% ±$\pm 1.65$
Social
Schimanski et al. (2023b) BERT 89.84% ±$\pm 0.79$
Schimanski et al. (2023b) RoBERTa 89.87% ±$\pm 1.35$
Schimanski et al. (2023b) DistilRoBERTa 90.59% ±$\pm 1.03$
Schimanski et al. (2023b) SocRoBERTa 91.90% ±$\pm 1.79$
Schimanski et al. (2023b) SocDistilRoBERTa 91.24% ±$\pm 1.86$
Governance
Schimanski et al. (2023b) BERT 76.86% ±$\pm 1.77$
Schimanski et al. (2023b) RoBERTa 77.03% ±$\pm 1.82$
Schimanski et al. (2023b) DistilRoBERTa 76.65% ±$\pm 2.39$
Schimanski et al. (2023b) GovRoBERTa 78.48% ±$\pm 2.62$
Schimanski et al. (2023b) GovDistilRoBERTa 78.86% ±$\pm 1.59$
(c) F1-scores (unspecified) on environment, social, and governance datasets (Schimanski et al., 2024a). F1-scores are reported as mean ±$\pm std$
Table 14. Reported performances for the datasets on ESG classification
Source Model Water Forest Biodiv. Nature
Schimanski et al. (2024a) EnvironmentalBERT 94.47% ±$\pm 1.37$ 95.37% ±$\pm 0.92$ 92.76% ±$\pm 1.01$ 94.19% ±$\pm 0.81$
Schimanski et al. (2024a) ClimateBERT 95.10% ±$\pm 1.13$ 95.34% ±$\pm 0.94$ 92.49% ±$\pm 1.03$ 93.50% ±$\pm 0.64$
Schimanski et al. (2024a) RoBERTa 94.55% ±$\pm 0.86$ 94.78% ±$\pm 0.48$ 92.46% ±$\pm 1.54$ 93.97% ±$\pm 0.26$
Schimanski et al. (2024a) DistilRoBERTa 94.98% ±$\pm 1.16$ 95.29% ±$\pm 0.65$ 92.29% ±$\pm 1.23$ 93.55% ±$\pm 0.72$
Schimanski et al. (2024a) Keywords - - 63.03% 61.00%
@@
(a) F1-scores (unspecified) on Nature-related topic classification (Schimanski et al., 2024a). F1-scores are reported as mean ±$\pm std$.
Source Model F1
Spokoyny et al. (2023) Majority 0.79%
Spokoyny et al. (2023) Random 5.05%
Spokoyny et al. (2023) SVM 48.02%
Spokoyny et al. (2023) BERT 54.74%
Spokoyny et al. (2023) RoBERTa 52.90%
Spokoyny et al. (2023) DistilRoBERTa 51.13%
Spokoyny et al. (2023) Longformer 54.79%
Spokoyny et al. (2023) SciBERT 51.83%
Spokoyny et al. (2023) ClimateBERT 52.97%
(b) F1-scores (macro average) on SciDCC (Mishra and Mittal, 2021).
Source Models F1
Spokoyny et al. (2023) Majority 3.65%
Spokoyny et al. (2023) Random 6.45%
Spokoyny et al. (2023) SVM 58.34%
Spokoyny et al. (2023) BERT 64.64%
Spokoyny et al. (2023) RoBERTa 65.22%
Spokoyny et al. (2023) DistilRoBERTa 63.61%
Spokoyny et al. (2023) Longformer 64.03%
Spokoyny et al. (2023) SciBERT 63.62%
Spokoyny et al. (2023) ClimateBERT 64.24%
(c) F1-scores (macro average) on the climaTOPIC (Spokoyny et al., 2023).
Source Model F1
Jain et al. (2023) all-mpnet-base-v2 43.7%
Jain et al. (2023) TF-IDF 69%
Jain et al. (2023) Word2Vec 72%
Jain et al. (2023) RoBERTa 87.19%
Jain et al. (2023) BERT 87.14%
Jain et al. (2023) ClimateBERT 85.20%
(d) F1-scores (weighted average) on ledger transaction classification (Jain et al., 2023).
Source Model F1
Vaid et al. (2022) FastText 63.8%
Vaid et al. (2022) BERT-Base 69.6%
Vaid et al. (2022) BERT-Large 69.5%
Vaid et al. (2022) RoBERTa-Base 73.4%
Vaid et al. (2022) RoBERTa-Large 73.5%
Vaid et al. (2022) DistilBERT 69.4%
Spokoyny et al. (2023) Majority 13.83%
Spokoyny et al. (2023) Random 16.71%
Spokoyny et al. (2023) SVM 51.81%
Spokoyny et al. (2023) BERT 71.78%
Spokoyny et al. (2023) RoBERTa 74.58%
Spokoyny et al. (2023) DistilRoBERTa 72.33%
Spokoyny et al. (2023) Longformer 72.28%
Spokoyny et al. (2023) SciBERT 70.50%
Spokoyny et al. (2023) ClimateBERT 71.83%
(e) F1-scores (macro average) on ClimateEng (Vaid et al., 2022).
@@
Table 15. Reported performances for the datasets on climate-related topic detection (other topics.)
C.3. In-depth Disclosure: Climate Risk Classification
@@
Friederich et al. (2021) SVM 29.0% 35.1% 20.4% 45.7% 59.9%
Friederich et al. (2021) distilBERT 44.4% 49.7% 24.1% 43.1% 55.8%
Friederich et al. (2021) RoBERTa 49.6% 44.6% 35.6% 52.8% 59.6%
@@
(a) F1-scores (macro average) for climate risk disclosure Friederich et al. (2021).
Source Model F1
Xiang and Fujii (2023) BERT-Base 93.1%
Xiang and Fujii (2023) BERT-Base (Domain-spe.) 95.5%
Xiang and Fujii (2023) TinyBERT 87.0%
Xiang and Fujii (2023) DistilBERT 89.9%
Xiang and Fujii (2023) ClimateBERT 87.5%
Xiang and Fujii (2023) CCLA+Max-Pooling 82.9%
Xiang and Fujii (2023) Bi-LSTM-Attention 71.9%
Xiang and Fujii (2023) POS-Bi-LSTM-Attention 88.2%
Xiang and Fujii (2023) DARE 89.4%
(b) F1-scores (macro average) on Risk/Opportunity classification (Xiang and Fujii, 2023).
Source Model F1
Webersinke et al. (2022) distilRoBERTa 82.5% $\pm$ 4.6
Webersinke et al. (2022) climateBERT 83.8% $\pm$ 3.6
Bingler et al. (2023) Naive Bayes 72%
Bingler et al. (2023) SVM + BoW 72%
Bingler et al. (2023) SVM + ELMo 75%
Bingler et al. (2023) climateBERT 80%
@@
(c) F1-scores (weighted average) on risk/opportunity detection Webersinke et al. (2022); Bingler et al. (2023). Webersinke et al. (2022)’s F1-scores are reported as mean ±$\pm std$. Bingler et al. (2023) also reported an uncertainty measure (p-value) which is not shown in this table for simplicity.
Source Model F1
Kölbel et al. (2021) Bag-of-Words 83.63%
Kölbel et al. (2021) TF-IDF 80.05%
Kölbel et al. (2021) BERT 94.66%
(d) F1-scores (unspecified) for risk classification (General/Transition risk/Physical Risk) (Kölbel et al., 2021).
@@
Table 16. Reported performances for the datasets on risk detection and classification.
C.4. Green Claim Detection
Source Model F1
Stammbach et al. (2023) Majority baseline 00.0%
Stammbach et al. (2023) Random baseline 33.5%
Stammbach et al. (2023) ClaimBuster RoBERTa 33.9%
Stammbach et al. (2023) Pledge Detection RoBERTa 26.4%
Stammbach et al. (2023) TF-IDF SVM 69.1%
Stammbach et al. (2023) Character n-gram SVM 70.9%
Stammbach et al. (2023) DistilBERT 83.7%
Stammbach et al. (2023) ClimateBERT 83.8%
Stammbach et al. (2023) RoBERTa_base 82.4%
Stammbach et al. (2023) RoBERTa_large 84.9%
(a) F1-scores (binary) on environmental claim detection (Stammbach et al., 2023).
Source Model F1
Woloszyn et al. (2021) Bin RoBERTa Bal 92.08%
Woloszyn et al. (2021) Bin BERTweet Bal 91.48%
Woloszyn et al. (2021) Bin Flair Bal 88.52%
Woloszyn et al. (2021) Bin RoBERTa Unbal 88.88%
Woloszyn et al. (2021) Bin BERTweet Unbal 84.30%
Woloszyn et al. (2021) Bin Flair Unbal 81.55%
@@ C.4. Green Claim Detection
(b) F1-scores (binary) on green claim detection (Woloszyn et al., 2021).
Table 17. Reported performances for the datasets on green claim detection.
C.5. Green Claim Characteristics
Source Model F1
Woloszyn et al. (2021) Mult RoBERTa Bal. 81.45%
Woloszyn et al. (2021) Mult BERTweet Bal. 75.14%
Woloszyn et al. (2021) Mult Flair Bal. 69.96%
Woloszyn et al. (2021) Mult RoBERTa Unbal. 76.82%
Woloszyn et al. (2021) Mult BERTweet Unbal. 59.71%
Woloszyn et al. (2021) Mult Flair Unbal. 38.15%
(a) F1-scores (unspecified) on Implicit/Explicit green claim detection (Woloszyn et al., 2021).
Source Model Accuracy
Schimanski et al. (2023a) ClimateBERT 96.2% ±$\pm 0.4$
Schimanski et al. (2023a) DistilRoBERTa 94.4% ±$\pm 0.7$
Schimanski et al. (2023a) RoBERTa-base 95.8% ±$\pm 0.6$
Schimanski et al. (2023a) GPT-3.5-turbo 92.0%
(b) Accuracy on net-zero/reduction classification (Schimanski et al., 2023a). Accuracy is reported as mean ±$\pm std$.
Source Model Commit. & Act. Specificity
Bingler et al. (2023) Naive Bayes 75% 75%
Bingler et al. (2023) SVM + BoW 76% 75%
Bingler et al. (2023) SVM + ELMo 79% 76%
Bingler et al. (2023) climateBERT 81% 77%
@@ C.5. Green Claim Characteristics
(c) F1-scores (weighted average) on Commitment&Action and Specificity classification (Bingler et al., 2023). Bingler et al. (2023) also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
@@
Table 18. Reported performances for the datasets on claim characteristics classification.
C.6. Green Stance Detection
Source Model F1
Vaghefi et al. (2022) GPT-2 67%
Vaghefi et al. (2022) climateGPT-2 72%
Claim
Leippold and Diggelmann (2020) ALBERT(FEVER) 32.85%
Xiang and Fujii (2023) BERT 80.7%
Xiang and Fujii (2023) RoBERTa 72.3%
Xiang and Fujii (2023) DistilRoBERTa 72.0%
Xiang and Fujii (2023) ClimateBERT 76.8%
Xiang and Fujii (2023) POS-Bi-LSTM-Att. 79.3%
Webersinke et al. (2022) DistilRoBERTa 74.8% ±$\pm 3.6$
Webersinke et al. (2022) ClimateBERT 75.7% ±$\pm 4.4$
Subset 95 examples
Wang et al. (2021) RoBERTa (FEVER) 49.2%
Wang et al. (2021) + ClimateFEVER 64.2%
Wang et al. (2021) + UDA ratio 18 71.8%
Evidence
Spokoyny et al. (2023) Majority 26.08%
Spokoyny et al. (2023) Random 30.62%
Spokoyny et al. (2023) BERT 62.47%
Spokoyny et al. (2023) RoBERTa 60.74%
Spokoyny et al. (2023) DistilRoBERTa 61.54%
Spokoyny et al. (2023) Longformer 60.82%
Spokoyny et al. (2023) SciBERT 62.68%
Spokoyny et al. (2023) ClimateBERT 61.54%
(a) F1-score (macro average) on ClimateFEVER Leippold and Diggelmann (2020). The prediction can be made at Claim level or for each Claim-Evidence pair (Evidence). Webersinke et al. (2022)’s F1-scores are reported as mean ±$\pm std$.
Source Model Acc. F1
Luo et al. (2020) Majority class 0.43 17%
Luo et al. (2020) Linear 0.62 60%
Luo et al. (2020) BERT 0.75 73%
Luo et al. (2020) Human 0.71 -
(b) Accuracy and F1-scores (macro average) on GWSD (Luo et al., 2020).
Source Model F1
Vaid et al. (2022) FastText 34.3%
Vaid et al. (2022) BERT-Base 46.4%
Vaid et al. (2022) BERT-Large 48.9%
Vaid et al. (2022) RoBERTa-Base 51.0%
Vaid et al. (2022) RoBERTa-Large 48.9%
Vaid et al. (2022) DistilBERT 44.8%
Spokoyny et al. (2023) Majority 29.68%
Spokoyny et al. (2023) Random 25.52%
Spokoyny et al. (2023) SVM 42.92%
Spokoyny et al. (2023) BERT 55.37%
Spokoyny et al. (2023) RoBERTa 59.69%
Spokoyny et al. (2023) DistilRoBERTa 52.51%
Spokoyny et al. (2023) Longformer 34.68%
Spokoyny et al. (2023) SciBERT 48.67%
Spokoyny et al. (2023) ClimateBERT 52.84%
(c) F1-scores (macro average) on ClimateStance (Vaid et al., 2022).
Source Model macro bin.
Baselines
Lai and Chen (2023) BERT - 58%
Lai and Chen (2023) RoBERTa - 89%
Lai and Chen (2023) ClimateBERT 90% 90%
SeqGAN
Lai and Chen (2023) BERT - 83%
Lai and Chen (2023) RoBERTa - 87%
Lai and Chen (2023) ClimateBERT 90% 90%
MaliGAN
Lai and Chen (2023) BERT - 74%
Lai and Chen (2023) RoBERTa - 88%
Lai and Chen (2023) ClimateBERT 91% 91%
RankGAN
Lai and Chen (2023) BERT - 82%
Lai and Chen (2023) RoBERTa - 87%
Lai and Chen (2023) ClimateBERT 91% 90%
(d) F1-scores (binary and macro average) on stance on climate change remediation efforts) (Lai and Chen, 2023).
Table 19. Reported performance for the datasets on stance detection (Part 1)

### v2

@@ Appendix C Reported Performances
(b) Performance on TCFD classification task [sampson_tcfd-nlp_nodate] with 11 labels
Source Model F1 (macro)
auzepy_evaluating_2023 BART+MNLI 56.68%
(c) Reported performance of the evaluation of the zero-shot approach for TCFD classification (into 11 categories) [auzepy_evaluating_2023]
Table 14: Reported performance for the datasets on TFCD classification
Source Model F1 (micro)
huangFinBERTLargeLanguage2020 FinBERT 89.6%
huangFinBERTLargeLanguage2020 BERT 86.9%
huangFinBERTLargeLanguage2020 NB 75.0%
huangFinBERTLargeLanguage2020 SVM 75.0%
huangFinBERTLargeLanguage2020 RF 78.3%
huangFinBERTLargeLanguage2020 CNN 81.9%
huangFinBERTLargeLanguage2020 LSTM 86.0%
(a) Performance for ESG classification [huangFinBERTLargeLanguage2020]
Source Model Acc Stdev
LEE2023119726 Multilingual BERT 84.09 % 0.092
LEE2023119726 KoBERT 85.98 % 0.163
LEE2023119726 DistilKoBERT 84.89 % 0.081
LEE2023119726 KB-ALBERT 85.62 % 0.026
LEE2023119726 KLUE-RoBERTa-base 86.05 % 0.125
LEE2023119726 KLUE-RoBERTa-large 86.66 % 0.099
(b) Performance on ESG classification with 4 labels LEE2023119726
Environment Models F1 (unspecified)
schimanski_bridging_2023 BERT 91.85% $\pm$ 1.25
schimanski_bridging_2023 RoBERTa 92.35% $\pm$ 2.29
schimanski_bridging_2023 DistilRoBERTa 90.97% $\pm$ 2.00
schimanski_bridging_2023 EnvRoBERTa 93.19% $\pm$ 1.40
schimanski_bridging_2023 EnvDistilRoBERTa 92.35% $\pm$ 1.65
Social Models
schimanski_bridging_2023 BERT 89.84% $\pm$ 0.79
schimanski_bridging_2023 RoBERTa 89.87% $\pm$ 1.35
schimanski_bridging_2023 DistilRoBERTa 90.59% $\pm$ 1.03
schimanski_bridging_2023 SocRoBERTa 91.90% $\pm$ 1.79
schimanski_bridging_2023 SocDistilRoBERTa 91.24% $\pm$ 1.86
Governance Models
schimanski_bridging_2023 BERT 76.86% $\pm$ 1.77
schimanski_bridging_2023 RoBERTa 77.03% $\pm$ 1.82
schimanski_bridging_2023 DistilRoBERTa 76.65% $\pm$ 2.39
schimanski_bridging_2023 GovRoBERTa 78.48% $\pm$ 2.62
schimanski_bridging_2023 GovDistilRoBERTa 78.86% $\pm$ 1.59
(c) F1-scores (mean $\pm$ std) for environment, social, and governance models under the standard hyperparameter setup.
Table 15: Reported performance for the datasets on ESG classification
Source Model F1 (macro)
spokoyny2023answering Majority 0.79%
spokoyny2023answering Random 5.05%
spokoyny2023answering SVM 48.02%
spokoyny2023answering BERT 54.74%
spokoyny2023answering RoBERTa 52.90%
spokoyny2023answering DistilRoBERTa 51.13%
spokoyny2023answering Longformer 54.79%
spokoyny2023answering SciBERT 51.83%
spokoyny2023answering ClimateBERT 52.97%
(a) Performance on SciDCC mishra2021neuralnere
Source Model F1 (weighted)
jain_supply_2023 all-mpnet-base-v2 43.7%
jain_supply_2023 TF-IDF 69%
jain_supply_2023 Word2Vec 72%
jain_supply_2023 RoBERTa 87.19%
jain_supply_2023 BERT 87.14%
jain_supply_2023 ClimateBERT 85.20%
(b) Ledger classification [jain_supply_2023]
Source Model Water (%) Forest (%) Biodiversity (%) Nature (%)
Schimanski2024nature EnvironmentalBERT 94.47 $\pm$ 1.37 95.37 $\pm$ 0.92 92.76 $\pm$ 1.01 94.19 $\pm$ 0.81
Schimanski2024nature ClimateBERT 95.10 $\pm$ 1.13 95.34 $\pm$ 0.94 92.49 $\pm$ 1.03 93.50 $\pm$ 0.64
Schimanski2024nature RoBERTa 94.55 $\pm$ 0.86 94.78 $\pm$ 0.48 92.46 $\pm$ 1.54 93.97 $\pm$ 0.26
Schimanski2024nature DistilRoBERTa 94.98 $\pm$ 1.16 95.29 $\pm$ 0.65 92.29 $\pm$ 1.23 93.55 $\pm$ 0.72
Schimanski2024nature Keywords - - 63.03 61.00
@@
(c) Performance (unspecified F1-score) on Nature related topic classification [Schimanski2024nature]
Source Models F1 (macro)
spokoyny2023answering Majority 3.65
spokoyny2023answering Random 6.45
spokoyny2023answering SVM 58.34
spokoyny2023answering BERT 64.64
spokoyny2023answering RoBERTa 65.22
spokoyny2023answering DistilRoBERTa 63.61
spokoyny2023answering Longformer 64.03
spokoyny2023answering SciBERT 63.62
spokoyny2023answering ClimateBERT 64.24
(d) Performance on the CLIMA-TOPIC classification [spokoyny2023answering]
Source Model F1 (macro)
vaid-etal-2022-towards FastText 63.8%
vaid-etal-2022-towards BERT-Base 69.6%
vaid-etal-2022-towards BERT-Large 69.5%
vaid-etal-2022-towards RoBERTa-Base 73.4%
vaid-etal-2022-towards RoBERTa-Large 73.5%
vaid-etal-2022-towards DistilBERT 69.4%
spokoyny2023answering Majority 13.83%
spokoyny2023answering Random 16.71%
spokoyny2023answering SVM 51.81%
spokoyny2023answering BERT 71.78%
spokoyny2023answering RoBERTa 74.58%
spokoyny2023answering DistilRoBERTa 72.33%
spokoyny2023answering Longformer 72.28%
spokoyny2023answering SciBERT 70.50%
spokoyny2023answering ClimateBERT 71.83%
(e) Results for the Fine-grained Classification using ClimateEng dataset.
@@
Table 16: Reported performance for the datasets on climate-related topic detection (other topics)
In-depth Disclosure: Climate Risk Classification
Source Model F1 (macro)
xiang_dare_2023 BERT-Base 93.1%
xiang_dare_2023 BERT-Base (Domain pretrained) 95.5%
xiang_dare_2023 TinyBERT 87.0%
xiang_dare_2023 DistilBERT 89.9%
xiang_dare_2023 ClimateBERT 87.5%
xiang_dare_2023 CCLA+Max-Pooling 82.9%
xiang_dare_2023 Bi-LSTM-Attention 71.9%
xiang_dare_2023 POS-Bi-LSTM-Attention 88.2%
xiang_dare_2023 DARE 89.4%
(a) Performance on sentiment analysis (Risk/Opportunity) [xiang_dare_2023]
Source Model F1 (unspecified)
kolbel_ask_2021 Bag-of-Words 83.63%
kolbel_ask_2021 TF-IDF 80.05%
kolbel_ask_2021 BERT 94.66%
(b) Performance for risk classification (General/Transition risk/Physical Risk) [kolbel_ask_2021]
@@
Friederich_climate_risk_disclosure SVM 29.0% 35.1% 20.4% 45.7% 59.9%
Friederich_climate_risk_disclosure distilBERT 44.4% 49.7% 24.1% 43.1% 55.8%
Friederich_climate_risk_disclosure RoBERTa 49.6% 44.6% 35.6% 52.8% 59.6%
@@
(c) Performance (F1 macro average) for climate risk disclosure Tasks from Friederich_climate_risk_disclosure
Source Model F1 (unspecified)
huangFinBERTLargeLanguage2020 FinBERT 87.8%
huangFinBERTLargeLanguage2020 BERT 84.2%
huangFinBERTLargeLanguage2020 NB 71.1%
huangFinBERTLargeLanguage2020 SVM 69.6%
huangFinBERTLargeLanguage2020 RF 66.8%
huangFinBERTLargeLanguage2020 CNN 72.5%
huangFinBERTLargeLanguage2020 LSTM 73.3%
(d) Performance on sentiment Analysis with FinBERT huangFinBERTLargeLanguage2020.
Source Model F1 (weighted average)
nicolas_webersinke_climatebert_2021 distilRoBERTa 82.5 % $\pm$ 4.6
nicolas_webersinke_climatebert_2021 climateBERT 83.8 % $\pm$ 3.6
bingler2023cheaptalkspecificitysentiment Naive Bayes 72 %
bingler2023cheaptalkspecificitysentiment SVM + BoW 72 %
bingler2023cheaptalkspecificitysentiment SVM + ELMo 75 %
bingler2023cheaptalkspecificitysentiment climateBERT 80 %
@@
(e) Performance on risk/opportunity detection task from nicolas_webersinke_climatebert_2021 and bingler2023cheaptalkspecificitysentiment. bingler2023cheaptalkspecificitysentiment also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
@@
Table 17: Reported performance for the datasets on sentiment analysis / risk detection
Green Claim Detection
Source Model F1 (binary)
stammbach_environmental_2023 Majority baseline 00.0%
stammbach_environmental_2023 Random baseline 33.5%
stammbach_environmental_2023 ClaimBuster RoBERTa 33.9%
stammbach_environmental_2023 Pledge Detection RoBERTa 26.4%
stammbach_environmental_2023 TF-IDF SVM 69.1%
stammbach_environmental_2023 Character n-gram SVM 70.9%
stammbach_environmental_2023 DistilBERT 83.7%
stammbach_environmental_2023 ClimateBERT 83.8%
stammbach_environmental_2023 RoBERTa_base 82.4%
stammbach_environmental_2023 RoBERTa_large 84.9%
(a) Performance on environmental claim detection from stammbach_environmental_2023
Source Model F1 (binary)
vinicius_woloszyn_towards_2021 Bin RoBERTa Bal 92.08%
vinicius_woloszyn_towards_2021 Bin BERTweet Bal 91.48%
vinicius_woloszyn_towards_2021 Bin Flair Bal 88.52%
vinicius_woloszyn_towards_2021 Bin RoBERTa Unbal 88.88%
vinicius_woloszyn_towards_2021 Bin BERTweet Unbal 84.30%
vinicius_woloszyn_towards_2021 Bin Flair Unbal 81.55%
@@
(b) Performance on green claim detection from vinicius_woloszyn_towards_2021
Table 18: Reported performance for the datasets on claim detection
Green Claim Characteristics
Source Model Commitments & Actions Specificity
F1 (weighted average) F1 (weighted average)
bingler2023cheaptalkspecificitysentiment Naive Bayes 75 % 75 %
bingler2023cheaptalkspecificitysentiment SVM + BoW 76 % 75 %
bingler2023cheaptalkspecificitysentiment SVM + ELMo 79 % * 76 %
bingler2023cheaptalkspecificitysentiment climateBERT 81 % 77 %
(a) Performance on downstream task from bingler2023cheaptalkspecificitysentiment.
@@
citetbingler2023cheaptalkspecificitysentiment also reported an uncertainty measure (p-value) which is not reported in this table for simplicity.
Source Model F1(unspecified)
vinicius_woloszyn_towards_2021 Mult RoBERTa Bal 81.45 %
vinicius_woloszyn_towards_2021 Mult BERTweet Bal 75.14 %
vinicius_woloszyn_towards_2021 Mult Flair Bal 69.96 %
vinicius_woloszyn_towards_2021 Mult RoBERTa Unbal 76.82 %
vinicius_woloszyn_towards_2021 Mult BERTweet Unbal 59.71 %
vinicius_woloszyn_towards_2021 Mult Flair Unbal 38.15 %
(b) Performance on Implicit/Explicit green claim detection from vinicius_woloszyn_towards_2021
Source Model Accuracy Std
tobias_schimanski_climatebert-netzero_2023 ClimateBERT 96.2 % 0.4
tobias_schimanski_climatebert-netzero_2023 DistilRoBERTa 94.4 % 0.7
tobias_schimanski_climatebert-netzero_2023 RoBERTa-base 95.8 % 0.6
tobias_schimanski_climatebert-netzero_2023 GPT-3.5-turbo 92.0 % -
(c) Performances on net-zero/reduction detection task from tobias_schimanski_climatebert-netzero_2023
@@
Table 19: Reported performance for the datasets on claim characteristics classification tasks
Green Stance Detection
Source Model ClimateFEVER
vaghefi2022deep GPT-2 67 %
vaghefi2022deep climateGPT-2 72 %
Claim F1 (macro)
diggelmann_climate-fever_2020 ALBERT(FEVER) 32.85%
xiang_dare_2023 BERT 80.7%
xiang_dare_2023 RoBERTa 72.3%
xiang_dare_2023 DistilRoBERTa 72.0%
xiang_dare_2023 ClimateBERT 76.8%
xiang_dare_2023 POS-Bi-LSTM-Attention 79.3%
nicolas_webersinke_climatebert_2021 DistilRoBERTa 74.8% $\pm$ 3.6
nicolas_webersinke_climatebert_2021 ClimateBERT 75.7% $\pm$ 4.4
subset 95 examples F1 (macro)
Wang2021EvidenceBA RoBERTa Large (FEVER) 49.2%
Wang2021EvidenceBA + CLIMATE-FEVER 64.2%
Wang2021EvidenceBA + UDA ratio 8 71.7%
Wang2021EvidenceBA + UDA ratio 18 71.8%
Evidence F1 (macro)
spokoyny2023answering Majority 26.08%
spokoyny2023answering Random 30.62%
spokoyny2023answering BERT 62.47%
spokoyny2023answering RoBERTa 60.74%
spokoyny2023answering DistilRoBERTa 61.54%
spokoyny2023answering Longformer 60.82%
spokoyny2023answering SciBERT 62.68%
spokoyny2023answering ClimateBERT 61.54%
(a) Performance on ClimateFEVER diggelmann_climate-fever_2020. The dataset contains pairs of claim and evidences, and an aggregated label for the claim. The evaluation is computed at the claim level (aggregated). Except for spokoyny2023answering, they computed the score at the evidence level (for each pair). $\pm$ value reports the standard deviation.
Source Model Acc F1 (macro)
luo_detecting_2020 Majority class 0.43 17%
luo_detecting_2020 Linear 0.62 60%
luo_detecting_2020 BERT 0.75 73%
luo_detecting_2020 Human 0.71 -
(b) Performance on GWSD[luo_detecting_2020] (stance detection on global warming)
Source Model F1 (macro)
vaid-etal-2022-towards FastText 34.3%
vaid-etal-2022-towards BERT-Base 46.4%
vaid-etal-2022-towards BERT-Large 48.9%
vaid-etal-2022-towards RoBERTa-Base 51.0%
vaid-etal-2022-towards RoBERTa-Large 48.9%
vaid-etal-2022-towards DistilBERT 44.8%
spokoyny2023answering Majority 29.68%
spokoyny2023answering Random 25.52%
spokoyny2023answering SVM 42.92%
spokoyny2023answering BERT 55.37%
spokoyny2023answering RoBERTa 59.69%
spokoyny2023answering DistilRoBERTa 52.51%
spokoyny2023answering Longformer 34.68%
spokoyny2023answering SciBERT 48.67%
spokoyny2023answering ClimateBERT 52.84%
(c) Performance on ClimateStance, the stance classification task of vaid-etal-2022-towards (3 labels: Favor, Against, Ambiguous)
Model F1 (macro) F1 (binary)
Baseline Models
BERT - 58%
RoBERTa - 89%
ClimateBERT 90% 90%
SeqGAN Approach
BERT - 83%
RoBERTa - 87%
ClimateBERT 90% 90%
MaliGAN Approach
BERT - 74%
RoBERTa - 88%
ClimateBERT 91% 91%
RankGAN Approach
BERT - 82%
RoBERTa - 87%
ClimateBERT 91% 90%
(d) Performance on stance on climate change remediation efforts (support/attack) [lai_using_2023]